The sum of a number and its square is 6. Find the number
Answer by scott8148(5891)

n + n^2 = 6

n^2 + n - 6 = 0

factoring ___ (n + 3)(n - 2) = 0

n + 3 = 0 ___ n = -3

n - 2 = 0 ___ n = 2

two solutions